OpenStudy (anonymous):

2 log7 -2r=0 Help please

OpenStudy (anonymous):

whats r??

OpenStudy (anonymous):

Variable

OpenStudy (michele_laino):

please divide by 2 both sides of your equation, first

OpenStudy (anonymous):

& then?

OpenStudy (solomonzelman):

\(\large\color{slate}{ 2\log(7)-2r=0 }\) \(\large\color{slate}{ 2\log(7)=2r }\) divide both sides by 2

OpenStudy (solomonzelman):

and remember that when the base is unspecified, it is \(\large\color{slate}{ 10 }\).